OpenStudy (anonymous):

A farmer along the boundary of a square field of side 10 min in 40 sec.What will be the magnitude of displacement of 2 min 20 sec ?

OpenStudy (anonymous):



OpenStudy (aaronq):

I dont think you pasted the question correctly.

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

@aryandecoolest the question is here

OpenStudy (anonymous):

 \[AC = \sqrt{(AB ^2 + BC ^2 )} = \sqrt{(100+100) } = 10√2 m\] REST OF ALL YOU KNOW :)

OpenStudy (anonymous):

He will complete 1 round in 40 s. so speed= 1m/sec So, distance covered in 2 min 20 s or 140 s is = 140 × 1 = 140 m Now, Number of rounds of the square completed in moving through 140 m is = 140/40 = 3.5 rounds so, In 3 rounds the displacement is zero. In 0.5 round the farmer reaches the diagonally opposite end of the square from his starting point Hope it's clear now :D @Greaty

OpenStudy (anonymous):

yes thank u @aryandecoolest u always help me.. :D

OpenStudy (anonymous):

oh!! glad to know, i try my best :) anytime ) @Greaty

OpenStudy (anonymous):

:D